# Conformal Temporal Logic Planning using Large Language Models

### v4

@@ II Problem Formulation
where ${\mathbf{p}}(t)$ stands for the state (e.g., position and orientation) of the robot, and ${\mathbf{u}}(t)$ stands for control input at discrete time $t$. We assume that the robot state ${\mathbf{p}}(t)$ is known for all time instants $t\geq 0$. The robot has $A>0$ number of abilities/skills collected in a set ${\mathcal{A}}\in\{1,\dots,A\}$. Each skill $a\in{\mathcal{A}}$ is represented as text such as ‘take a picture’, ‘grab’, or ‘move to’. Application of a skill $a$ at an object/region with location ${\mathbf{x}}$ at time $t\geq 0$ is denoted by $s(a,{\mathbf{x}},t)$ or, for brevity, when it is clear from the context, by $s(t)$. The time step $t$ is increased by one, once an action is completed. We assume that the robot has access to low level controllers to apply the skills in ${\mathcal{A}}$.
@@
Partially Known Semantic Environment: The robot operates within a semantic environment $\Omega\subseteq\mathbb{R}^{d}$, $d\in\{2,3\}$ with fixed, static, and potentially unknown obstacle-free space denoted by $\Omega_{\text{free}}\subseteq\Omega$. The space $\Omega_{\text{free}}$ is populated with $M>0$ static semantic objects. Each object $e$ is characterized by its location ${\mathbf{x}}_{e}$ and semantic label $o_{e}\in{\mathcal{O}}$, where ${\mathcal{O}}$ is a set collecting all possible semantic labels that objects may have (e.g., ‘bottle’ or ‘chair’). The robot is assumed to have knowledge of both the location and label of each object. Objects may also be located inside containers (e.g., drawer or fridge), with their status (open/closed) initially unknown. Also, the occupied space $\Omega\setminus\Omega_{\text{free}}$ may prevent access to certain semantic objects rendering execution of certain decisions $s(t)$ infeasible. We assume that the robot is equipped with sensors allowing it to detect obstacles and reason about the status of containers containing objects of interest. Hereafter, we denote by ${\mathcal{S}}$ a finite set collecting all decisions $s(t)$ the robot can take. This set is constructed offline using the action space ${\mathcal{A}}$ and the available objects where each action can be applied.
@@
Mission Specification: The robot is tasked with a high-level mission with temporal and logical requirements. To formally define such tasks, we employ Linear Temporal Logic (LTL). LTL is a formal language that comprises a set of atomic propositions (AP) (i.e., Boolean variables), denoted by $\mathcal{AP}$, Boolean operators, (i.e., conjunction $\wedge$, and negation $\neg$), and temporal operators, such as always $\square$, eventually $\lozenge$, and until $\mathcal{U}$; see also Example II.4. A formal presentation of the syntax and semantics of LTL can be found in [82]. For simplicity, hereafter, we restrict our attention to co-safe LTL formulas that is a fragment of LTL that exclude the ‘always’ operator. Co-safe LTL formulas can be satisfied within a finite horizon $H$. For brevity, throughout the paper, we may use the term LTL instead of co-safe LTL. We define APs so that they are true when a sub-task expressed in natural language (NL) is satisfied, and false otherwise. Valid definitions of APs include sub-tasks that can be accomplished by a finite robot trajectory $\tau$, defined as a finite sequence of $T$ decisions selected from ${\mathcal{S}}$, i.e.,
$\tau=s(t),s(t+1),\dots,s(t+k),\dots,s(t+T-1),$ (2)
@@
where $k\in\{0,\dots,T-1\}$, for some $T\geq 1$; see also Example II.4 and Section V-D. We call formulas constructed in this way as co-safe LTL-NL formulas. Co-safe LTL-NL formulas are satisfied by finite-horizon robot trajectories $\tau_{\phi}$ defined as
$\tau_{\phi}=\tau_{1},\dots,\tau_{n},\dots,\tau_{N},$ (3)
@@
Our goal is to design plans $\tau_{\phi}$ satisfying co-safe LTL-NL tasks under the following assumptions:
Assumption II.1 (Error-free Skills)
The robot has access to low-level controllers allowing it to apply skills error-free.
Assumption II.2 (Independent APs)
The APs are independent of each other, meaning that the satisfaction of one AP does not prohibit the satisfaction of another AP.
Remark II.3 (Assumptions)
We make Assumption II.1 as we consider robots with known and deterministic system dynamics as in (1). This assumption will be used to provide mission completion guarantees in Section IV; see Theorem IV.2. Formally relaxing this assumption and extending the proposed planner to account for uncertainty in the system dynamics and the performance of the robot skills is part of our future work; see also Remark IV.3. Assumption II.2 is made only for simplicity as it will be discussed in Remark III.1.
Problem Statement: This paper addresses the following problem (see Ex. II.4):
Problem 1
Given a robot with capabilities ${\mathcal{A}}$, a partially unknown semantic environment $\Omega$, and a co-safe LTL-NL task $\phi$, design a robot plan $\tau_{\phi}$ satisfying $\phi$.
Example II.4
@@
III Hierarchical Temporal Logic Planning with Natural Language Instructions
In this section, we propose HERACLEs, a hierarchical planner to address Problem 1. In Section III-A, we present an overview HERACLEs; see Alg. 1. A detailed description of its components is provided in Sections III-B-III-E.
@@ III Hierarchical Temporal Logic Planning with Natural Language Instructions
2:Symbolic planner generates sub-task: $\pi_{\text{next}}$ and $\Sigma^{\text{unsafe}}$
3:Convert sub-task into an initial prompt $\ell(1)$ with empty history of action
@@
7:         Compute the prediction set $\mathcal{C}(\ell(t+k))$ (Alg. 2)
8:         if  $|\mathcal{C}(\ell(t+k))|>1$  then
@@
9:              Request a new sub-task ($\pi_{\text{next}}$, $\Sigma^{\text{unsafe}}$)
10:              if $\pi_{\text{next}}\neq\varnothing$ then
11:                  $t=t+k$, and go to line 5
12:              else
13:                  Obtain $s(t+k)$ from human operator
14:         else
15:              Pick (unique) decision $s(t+k)\in\mathcal{C}(\ell(t+k))$
@@
16:         if $\Omega$ prevents execution of $s(t+k)$ then
@@
17:              Request a new sub-task ($\pi_{\text{next}}$, $\Sigma^{\text{unsafe}}$)
18:              if $\pi_{\text{next}}\neq\varnothing$ then
19:                  $t=t+k$, and go to line 5
20:              else
21:                  Report mission failure (termination)
22:         else
23:              Execute decision $s(t+k)$
24:              Observe environment update $p(t+k)$
25:         Update $\ell(t+k+1)=\ell(t+k)+s(t+k)+p(t+k)$
26:     Construct $\tau=[s(t),\dots,s(t+T-1)]$
27:     Append $\tau$ to the plan $\tau_{\phi}$
28:     Current time step: $t=t+T$
@@
29:     Request a new sub-task ($\pi_{\text{next}}$, $\Sigma^{\text{unsafe}}$) and check for mission accomplishment
III-A Overview of HERACLEs
@@
HERACLEs takes as input an LTL-NL mission defined over a partially known semantic environment. This LTL-NL formula is processed online by an existing symbolic temporal logic planner that, given the current mission status, determines the next language-based sub-task the robot should accomplish to make mission progress. This symbolic planner is presented in Section III-B and is adopted from [19, 20]; we note that any other temporal logic planner can be used. This language-based sub-task serves as an input to a pre-trained LLM that is responsible for generating a feasible plan of the form (2); see Section III-C. The LLM planner communicates back to the LTL planner with one of the following messages: (i) The LLM is ‘uncertain’ about what the correct plan is [line 7-8, Alg. 1]. In this case, the LTL planner generates an alternative sub-task to proceed [lines 9-11, Alg. 1]. If such a sub-task does not exist, help from users is requested who either provide the correct decision or decide to halt operation [line 13, Alg. 1]. (ii) A plan has been designed but the (initially unknown) geometric structure of the environment ends up preventing the robot from fully executing it [line 16, Alg. 1]. In this case, the LTL planner generates an alternative sub-task to proceed and the above process repeats [lines 17-19, Alg. 1]. If such a sub-task does not exist, the mission terminates unsuccessfully (e.g., the LTL-NL mission may be infeasible) [line 21, Alg. 1]. (iii) A plan has been designed and executed, accomplishing the assigned sub-task [lines 22-24, Alg. 1]. In this case, the LTL planner generates a new sub-task, if the overall mission has not been accomplished yet, and the above process repeats [line 29, Alg. 1]. A detailed description of when the LLM requests assistance (cases (i)-(ii)) from the symbolic planner is provided in Section III-D. Conformal prediction, outlined in Section III-E, is employed for uncertainty quantification in case (i).
III-B Symbolic Temporal Logic Task Planner: What Language-based Sub-Task to Accomplish Next?
@@
In this section, we provide an overview of an existing temporal logic task planner that we employ to determine what language sub-task the robot should accomplish next to make progress towards accomplishing the LTL-NL task $\phi$. A detailed presentation of this planner can be found in [19, 20].
@@
Consider an LTL-NL specification $\phi$. Initially, we translate $\phi$ into a Deterministic Finite state Automaton (DFA). This step occurs offline. The DFA can be conceptualized as a graph structure, with state-space (set of nodes) and transitions (set of edges) between states/nodes. The DFA has an initial node representing the mission’s commencement, a set of final nodes denoting mission completion, and intermediate nodes representing various stages of the mission. Transitions between nodes occur when specific atomic predicates are satisfied. Mission accomplishment is achieved upon reaching a final state starting from the initial one.
At any time $t\geq 0$, the task planner receives three inputs: (i) an LTL-NL task $\phi$; (ii) the current mission status; and (iii) a set $\Sigma^{\text{unc}}(t)$ of NL-based APs/sub-tasks. The input (ii) is represented by the DFA state that has been reached, starting from the initial one, given the sequence of actions that the robot has applied up to time $t$. The set in (iii) is initially empty, i.e., $\Sigma^{\text{unc}}(0)=\emptyset$, and is dynamically updated as the robot navigates the world. Specifically, it is updated so that it collects all NL-based APs that the robot is uncertain about how to accomplish or that are physically impossible to achieve (e.g., they require reaching inaccessible objects). The details of its update are discussed in Section III-C.
@@
Given these inputs, the task planner generates: (i) an NL-based AP, denoted by $\pi_{\text{next}}$; and (ii) a set of NL-based APs collected in a set $\Sigma^{\text{unsafe}}$ [line 2, Alg. 1]. These sets impose two conditions that must be met to make mission progress (i.e., to move closer to a final DFA node). Specifically, the robot should design a plan $\tau$, defined as in (2), to (a) accomplish the language-based sub-task captured in $\pi_{\text{next}}$ while, (b) in the meantime, not satisfying any of the APs collected in $\Sigma^{\text{unsafe}}$. Note that the set $\Sigma^{\text{unsafe}}$ depends on the current mission status and, therefore, it does not remain fixed throughout the mission. These two requirements act as inputs to an LLM-based planner; refer to Section III-C. The LLM-based planner is responsible for designing a plan $\tau$ that satisfies (a)-(b).
Remark III.1 (Relaxing Assumption II.2)
@@
The temporal logic planner used in Section III-B is myopic as it computes the next NL-based AP $\pi_{\text{next}}$ for the robot, instead of a sequence of sub-tasks that need to be accomplished to satisfy the overall mission. Assumption II.2 is a necessary condition required to ensure completeness of that symbolic planner. This assumption can be relaxed by employing non-myopic symbolic planners, which generate a sequence of predicates/sub-tasks (rather than only the next sub-task) that the robot should accomplish to complete the mission [14, 28]. The benefit of using the employed myopic symbolic planner, as opposed to a non-myopic one, lies in its computational efficiency, as demonstrated in [19, 20].
III-C LLM-based Planner: How to Accomplish the Assigned Language-based Sub-Task?

### v5

@@ 2. Problem Formulation
Robot System and Skills: Consider a robot governed by the following deterministic and known dynamics: ${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t))$, where ${\mathbf{p}}(t)$ stands for the robot state (e.g., position and orientation), and ${\mathbf{u}}(t)$ stands for control input at discrete time $t$. The robot has $A>0$ number of abilities/skills collected in a set ${\mathcal{A}}\in\{1,\dots,A\}$. Each skill $a\in{\mathcal{A}}$ is represented as text such as ‘take a picture’, ‘grab’, or ‘move to’. Application of a skill $a$ at an object/region with location $\mathbf{x}$} at time $t\geq 0$ is denoted by $s(a,{\mathbf{x}},t)$ or, for brevity, when it is clear from the context, by $s(t)$. The time step $t$ is increased by one, once an action is completed. We assume that the robot has access to low-level controllers to apply the skills in $\mathcal{A}$}.
@@
Partially Known Semantic Environment: The robot operates within a semantic environment $\Omega\subseteq\mathbb{R}^{d}$, $d\in\{2,3\}$ with fixed, static, and potentially unknown obstacle-free space denoted by $\Omega_{\text{free}}\subseteq\Omega$. The space $\Omega_{\text{free}}$ is populated with $M>0$ static semantic objects. Each object $e$ is characterized by its location ${\mathbf{x}}_{e}$ and semantic label $o_{e}$ (e.g., ‘bottle’ or ‘chair’). The robot is assumed to have knowledge of both the location and label of each object. Objects may also be located inside containers (e.g., drawer or fridge), with their status (open/closed) initially unknown. Also, the occupied space $\Omega\setminus\Omega_{\text{free}}$ may prevent access to certain semantic objects rendering execution of certain decisions $s(t)$ infeasible. We assume that the robot is equipped with sensors allowing it to detect obstacles and reason about the status of containers containing objects of interest. Hereafter, we denote by $\mathcal{S}$} a finite set collecting all decisions $s(t)$ the robot can take. This set is constructed offline using the action space $\mathcal{A}$} and the corresponding available objects.
@@
Mission Specification: The robot is tasked with a high-level mission with temporal and logical requirements. To formally define such tasks, we employ Linear Temporal Logic (LTL). LTL is a formal language that comprises a set of atomic propositions (AP) (i.e., Boolean variables), denoted by $\mathcal{AP}$, Boolean operators, (i.e., conjunction $\wedge$, and negation $\neg$), and temporal operators, such as always $\square$, eventually $\lozenge$, and until $\mathcal{U}$; see also Example 2.3. A formal presentation of the syntax and semantics of LTL can be found in (Baier and Katoen, 2008). For simplicity, hereafter, we restrict our attention to co-safe LTL formulas that are a fragment of LTL and that exclude the ‘always’ operator. Such formulas can be satisfied within a finite horizon $H$. Throughout the paper, for brevity, we will use the term LTL instead of co-safe LTL. We define APs so that they are true when a sub-task expressed in natural language (NL) is satisfied, and false otherwise. Valid definitions of APs include sub-tasks that can be accomplished by a finite robot trajectory $\tau$, defined as a finite sequence of $T$ decisions selected from $\mathcal{S}$}, i.e.,
(1) $\tau=s(t),s(t+1),\dots,s(t+k),\dots,s(t+T-1),$
@@
where $k\in\{0,\dots,T-1\}$, for some $T\geq 1$; see also Example 2.3 and Section 5.4. We call formulas constructed in this way as co-safe LTL-NL formulas. Co-safe LTL-NL formulas are satisfied by finite-horizon robot trajectories $\tau_{\phi}$ defined as
(2) $\tau_{\phi}=\tau_{1},\dots,\tau_{n},\dots,\tau_{N},$
@@
Distribution of Mission Scenarios: We formalize the above by defining a mission scenario as $\xi_{i}=\{{\mathcal{A}}_{i},\phi_{i},H_{i},\Omega_{i}\}$. Recall that ${\mathcal{A}}_{i},\phi_{i},H_{i},\text{~and~}\Omega_{i}$ refer to the robot skills, the LTL-NL mission, the mission horizon, and the semantic environment, respectively. The subscript $i$ is used to emphasize that these parameters can vary across scenarios. When it is clear from context, we drop the dependence on $i$. We assume that all scenarios are sampled from a distribution $\mathcal{D}$} defined over feasible scenarios $\xi$. Note that $\mathcal{D}$} is unknown, but we assume that we can sample i.i.d. scenarios from it as e.g., in (Ren et al., 2023).
Problem Statement: Our goal is to design safe LTL-NL planners, i.e., planners that can generate correct plans $\tau_{\phi_{i}}$ satisfying mission scenarios $\xi_{i}$ in at least $(1-\alpha)\%$ of the scenarios drawn from $\mathcal{D}$} for a user-specified $\alpha\in(0,1)$ under the following assumption (see Rem. 2.2):
Assumption 2.1 (Error-free Skills).
The robot has access to low-level controllers, allowing it to apply its skills error-free; see Appendix C for a discussion on how this assumption can be relaxed.
Remark 2.2 (Error-Free Skills & Probabilistic Task Satisfaction).
To address this problem, we will employ LLMs to process the NL-based predicates. The probabilistic satisfaction of a mission arises exclusively due to imperfections of the employed LLM and not due to low-level uncertainties in the execution of robot skills. The latter is precluded by Assumption 2.1. We note that Assumption 2.1 is common in related temporal logic (e.g., (Kantaros and Zavlanos, 2020; Kamale et al., 2023)) and language-based (e.g., (Chen et al., 2023b; Ren et al., 2023)) planners that also aim to design semantically correct plans that, when executed properly, fulfill the task requirements. Formally relaxing this assumption and extending the proposed planner to account for uncertainty in the system dynamics and the performance of the robot skills is part of our future work; see Appendix C.
Example 2.3 (LTL-NL Mission).
@@
3. Proposed Solution: HiERArchical ConformaL natural languagE planner (HERACLEs)
@@
2:Initialize $\Sigma_{\text{unc}}(0)=\emptyset$; LTL planner generates NL-based sub-task ($\pi_{\text{next}}$, $\Sigma_{\text{unsafe}}$)
3:Convert sub-task into an initial prompt $\ell(1)$ with empty history of actions
@@
7:   Compute prediction set $\mathcal{C}(\ell(t+k))$ (Alg. 2)
8:   if  $|\mathcal{C}(\ell(t+k))|>1$  then $\triangleright$ Ask-for-Help: Non-Singleton Prediction Set
9:     Ask human operator to choose decision from $\mathcal{C}(\ell(t+k))$
10:     if Human operator halts operation then $\triangleright$ No correct action in ${\mathcal{C}}({\ell}(t+k))$
11:      Update $\Sigma_{\text{unc}}$: $\Sigma_{\text{unc}}(t+k)\leftarrow\Sigma_{\text{unc}}(t+k)\cup\{\pi_{\text{next}}\}$ (if $\pi_{\text{next}}$ is not already in $\Sigma_{\text{unc}}(t+k)$)
@@
12:      Request new NL-based sub-task ($\pi_{\text{next}}$, $\Sigma_{\text{unsafe}}$) from LTL planner
13:      if $\pi_{\text{next}}\neq\varnothing$ then
14:       $t\leftarrow t+k$, and go to line 5
15:      else
16:       Report mission failure (termination)
17:     else
18:      Obtain $s(t+k)\in\mathcal{C}(\ell(t+k))$ from human operator
19:   else
20:     Pick (unique) decision $s(t+k)\in\mathcal{C}(\ell(t+k))$
@@
21:   if $\Omega$ prevents execution of $s(t+k)$ then $\triangleright$ Ask-for-Help: Physical Environment
22:     Update $\Sigma_{\text{unc}}$: $\Sigma_{\text{unc}}(t+k)\leftarrow\Sigma_{\text{unc}}(t+k)\cup\{\pi_{\text{next}}\}$ (if $\pi_{\text{next}}$ is not already in $\Sigma_{\text{unc}}(t+k)$)
@@
23:     Request new NL-based sub-task ($\pi_{\text{next}}$, $\Sigma_{\text{unsafe}}$) from LTL planner
24:     if $\pi_{\text{next}}\neq\varnothing$ then
25:      $t\leftarrow t+k$, and go to line 5
26:     else
27:      Report mission failure (termination)
28:   else
29:     Execute decision $s(t+k)$
30:     Observe environment update $p(t+k)$
31:   Update $\ell(t+k+1)\leftarrow\ell(t+k)+s(t+k)+p(t+k)$
32:  Remove all subtasks from $\Sigma_{\text{unc}}$ that were included because of LLM uncertainty (i.e., in line 11).
33:  Construct $\tau=[s(t),\dots,s(t+T-1)]$
34:  Append $\tau$ to the plan $\tau_{\phi}$
35:  Current time step: $t\leftarrow t+T$
@@
36:  Request new NL-based sub-task ($\pi_{\text{next}}$, $\Sigma_{\text{unsafe}}$) from LTL planner and check for mission accomplishment
In this section, we propose HERACLEs, a hierarchical planner to address the considered LTL-NL planning problem. In Section 3.1, we present an overview of HERACLEs; see Alg. 1. A detailed description of its components is provided in Sections 3.2-3.5.
3.1. Overview of HERACLEs
@@
HERACLEs takes as input an LTL-NL mission defined over a partially known semantic environment. This LTL-NL formula is processed online by an existing symbolic temporal logic planner that, given the current mission status, determines the next language-based sub-task the robot should accomplish to make mission progress. This symbolic planner is presented in Section 3.2 and is adopted from (Kantaros et al., 2020; Vasilopoulos et al., 2021); we note that any other temporal logic planner can be used. This language-based sub-task serves as an input to a pre-trained LLM that is responsible for generating a feasible plan of the form (1); see Section 3.3. The LLM planner communicates back to the LTL planner with one of the following messages: (i) The LLM is ‘uncertain’ about what the correct plan is [lines 7-8, Alg. 1]. In this case, help is requested from the human operator, who either provides the correct decision or chooses to halt the operation [lines 918, Alg.1]. If the operator decides to halt the operation, the LTL planner attempts to generate an alternative sub-task to continue the mission [lines 1114, Alg.1]. If no such sub-task exists, the mission terminates unsuccessfully (e.g., the LTL-NL mission may be infeasible) [line 16, Alg. 1]. (ii) A plan has been designed but the (initially unknown) geometric structure of the environment ends up preventing the robot from fully executing it [line 21, Alg. 1]. In this case, the LTL planner generates an alternative sub-task to proceed, and the above process repeats [lines 22-25, Alg. 1]. If such a sub-task does not exist, the mission terminates unsuccessfully [line 27, Alg. 1]. (iii) A plan has been designed and executed, accomplishing the assigned sub-task [lines 28-30, Alg. 1]. In this case, the LTL planner generates a new sub-task if the overall mission has not been accomplished yet, and the above process repeats [line 36, Alg. 1]. A detailed description of when the LLM requests assistance (cases (i)-(ii)) from the symbolic planner is provided in Section 3.4. Conformal prediction, outlined in Section 3.5, is employed for uncertainty quantification in case (i).
3.2. Symbolic Task Planner: What NL-based Sub-Task to Accomplish Next?
@@
In this section, we provide an overview of a symbolic/LTL task planner, presented in (Kantaros et al., 2020; Vasilopoulos et al., 2021), that we employ to determine what language sub-task the robot should accomplish next to make progress towards accomplishing the LTL-NL task $\phi$. Its input-output structure is illustrated in Fig. 2; see also Ex. 3.1. We emphasize that any other temporal logic planner can be used as long as it is complete, i.e., if there exists a sub-task to make progress, the employed planner will compute it.
Figure 2. Input-output structure of the symbolic planner. The output serves as an input to the neural planner (Fig. 3).
@@
Consider a mission scenario drawn from $\mathcal{D}$} associated with an LTL-NL specification $\phi$. Initially, we translate $\phi$ into a Deterministic Finite state Automaton (DFA) (Baier and Katoen, 2008); this step is performed offline, before robot deployment. The DFA can be viewed as a directed graph, with a set of nodes/states and edges/transitions between states. It has an initial node representing the mission’s commencement, a set of final nodes denoting mission completion, and intermediate nodes representing various stages of the mission. Transitions between nodes occur when specific atomic predicates are satisfied. Mission accomplishment is achieved upon reaching a final node while starting from the initial node.
@@
The symbolic task planner receives three inputs: (i) an LTL-NL task $\phi$; (ii) the current mission status at time $t$; and (iii) a set $\Sigma_{\text{unc}}(t)$ of NL-based APs. The input (ii) is represented by the DFA state that has been reached, starting from the initial one, given the sequence of actions that the robot has applied up to the current time step $t$. Initially, $\Sigma_{\text{unc}}(0)=\emptyset$ [line 2, Alg. 1]. As the robot navigates, $\Sigma_{\text{unc}}$ collects each NL-based AP for which either the LLM is uncertain about how to design a feasible plan, or whose satisfaction is physically impossible due to the (partially unknown) geometry of the environment. The details of its update are discussed in Section 3.4. Given these inputs, the symbolic task planner generates: (i) an NL-based AP, denoted by $\pi_{\text{next}}$; and (ii) a set of NL-based APs collected in a set $\Sigma_{\text{unsafe}}$ [line 2, Alg. 1]. These sets impose two conditions that must be met to make mission progress (i.e., to move closer to a final DFA node). Specifically, the robot should design a plan $\tau$, defined as in (1), to (a) accomplish the NL-based sub-task captured in $\pi_{\text{next}}$ while, (b) in the meantime, not satisfying any of the APs collected in $\Sigma_{\text{unsafe}}$. Note that the set $\Sigma_{\text{unsafe}}$ depends on the current mission status and, therefore, it does not remain fixed throughout the mission. These two requirements act as inputs to an LLM-based planner that is responsible for designing a plan $\tau$ that satisfies (a)-(b); see Section 3.3.
Figure 3. Overview of the neural planner. Given an NL-encoded sub-task from the symbolic planner, the neural planner generates robot actions and requests help—from either the symbolic planner or the user (see Fig. 5)—in cases of high uncertainty or infeasible sub-tasks. Once the sub-task is completed, the mission status and the set $\Sigma_{\text{unc}}$ are updated back to the symbolic planner.
@@
Notice that the temporal logic planner discussed above is myopic in the sense that it computes the next NL-based sub-task $\pi_{\text{next}}$ for the robot, instead of a sequence of sub-tasks that need to be accomplished to satisfy the overall mission. A necessary condition required to ensure completeness of that symbolic planner is that the APs are independent of each other, which means that the satisfaction of one AP does not prohibit the satisfaction of another AP. This assumption can be relaxed by employing non-myopic symbolic planners, which generate a sequence of sub-tasks (rather than only the next sub-task) that the robot should accomplish to complete the mission (Kantaros and Zavlanos, 2020; Shoukry et al., 2017). The benefit of using the employed myopic symbolic planner, as opposed to a non-myopic one, lies in its computational efficiency, as demonstrated in (Kantaros et al., 2020; Vasilopoulos et al., 2021).
Example 3.1 (Symbolic Planner).
Consider the LTL-NL formula $\phi$ discussed in Example 2.3. At time $t=0$, the symbolic planner generates the sub-task specified by $\pi_{\text{next}}=\pi_{1}$ and sets $\Sigma_{\text{unsafe}}=\{\pi_{4}\}$.
3.3. LLM-based Action Planner: How to Accomplish the Assigned NL-based Sub-Task?